# Cuspidal edges on focal surfaces of regular surfaces with parabolic points

### v1

@@ 4.1. Geometric invariants at cuspidal edges of focal surfaces
by (2.6), where $\gamma(v)$ is a singular curve for $C_{1}$ with $\gamma(v_{0})=p$ and $\hat{\gamma}=C_{1}\circ\gamma$. By (2.1), (4.5), (4.8) and Lemma 2.1, we have
$|\hat{\gamma}^{\prime}(v_{0})|^{2}$ $=\frac{1}{\kappa_{1}(p)^{4}}\left(\kappa_{1}(p)^{4}G(p)+(\kappa_{1})_{v}(p)^{2}\right),$
$\left\langle{\hat{\gamma}^{\prime\prime}(v_{0})},{\bm{e}_{1}(p)}\right\rangle$ $=\Gamma^{1}_{22}(p)\sqrt{E(p)}=-\frac{G_{u}}{2\sqrt{E(p)}}=-\frac{(\kappa_{2})_{u}(p)G(p)}{\kappa_{1}(p)\sqrt{E(p)}}.$
Thus we obtain the assertion. ∎
Corollary 4.4.
With the same situations in Proposition 4.3, the point $p$ is a $\partial_{u}$-subparabolic point for $f$ if and only if $\kappa_{\nu}^{C_{1}}$ vanishes at $p$. Moreover, $\bm{e}_{1}$ is singular at $p$ when $\kappa_{\nu}^{C_{1}}(p)=0$. Furthermore, the Gaussian curvature $K^{C_{1}}$ of $C_{1}$ is rationally bounded at $p$ in that case.
Proof.
The first assertion follows from Proposition 4.3. We get the second statement by Lemma 2.3. The last assertion holds by (A.6) and [16, Corollary C]. ∎
For the singular curvature $\kappa_{s}^{C_{1}}$ of $C_{1}$, we have the following.
Proposition 4.5.
Let $f\colon U\to\mathbb{R}^{3}$ be a regular surface parametrized by curvature line coordinates $(u,v)$. Suppose that $\kappa_{2}(p)=0$ and the focal surface $C_{1}$ is a cuspidal edge at $p$. Then the singular curvature $\kappa_{s}^{C_{1}}$ of $C_{1}$ is
$\kappa_{s}^{C_{1}}=\operatorname{sgn}((\kappa_{1})_{uu})\frac{\kappa_{1}^{3}\sqrt{G}}{(\kappa_{1})_{uu}\delta^{3}}\left(\kappa_{1}\det\operatorname{Hess}(\kappa_{1})-(\kappa_{1})_{v}(\kappa_{1})_{uu}\left(\Gamma^{2}_{22}\kappa_{1}+2(\kappa_{1})_{v}-(\kappa_{2})_{v}\right)\right)$
at $p$, where we set
$\delta=\sqrt{\kappa_{1}(p)^{4}G(p)+(\kappa_{1})_{v}(p)^{2}}.$
Proof.
Since $\bm{e}_{1}$ is a unit normal vector to $C_{1}$, the singular curvature $\kappa_{s}^{C_{1}}$ at $p$ is
$\kappa_{s}^{C_{1}}(p)=\operatorname{sgn}(\det(\gamma^{\prime},\eta)(v_{0})\cdot\eta^{C_{1}}\lambda^{C_{1}}(p))\frac{\det(\hat{\gamma}^{\prime}(v_{0}),\hat{\gamma}^{\prime\prime}(v_{0}),\bm{e}_{1}(p))}{|\hat{\gamma}^{\prime}(v_{0})|^{3}}.$
@@
$\bm{e}_{1}(p)\times\hat{\gamma}^{\prime}(v_{0})=\frac{(\kappa_{1})_{v}(p)}{\kappa_{1}(p)^{2}}\bm{e}_{2}(p)+|f_{v}(p)|\nu(p).$
@@
$\det(\hat{\gamma}^{\prime}(v_{0}),\hat{\gamma}^{\prime\prime}(v_{0}),\bm{e}_{1}(p))$
$=\frac{-|f_{v}(p)|}{\kappa_{1}(p)^{3}(\kappa_{1})_{uu}(p)}\Big{(}\kappa_{1}(p)\det\operatorname{Hess}(\kappa_{1})(p)$
$\qquad\qquad-(\kappa_{1})_{v}(p)(\kappa_{1})_{uu}(p)\left(\kappa_{1}(p)\Gamma^{2}_{22}(p)+2(\kappa_{1})_{v}(p)-(\kappa_{2})_{v}(p)\right)\Big{)}.$
@@
$\eta^{C_{1}}\lambda^{C_{1}}(p)=\frac{(\kappa_{1})_{uu}(p)}{\kappa_{1}(p)^{2}}|f_{v}(p)|,\quad\det(\gamma^{\prime},\eta)(v_{0})=-1$
@@
$\operatorname{sgn}(\det(\gamma^{\prime},\eta)(v_{0})\cdot\eta^{C_{1}}\lambda^{C_{1}}(p))=-\operatorname{sgn}((\kappa_{1})_{uu}(p)).$
Since
$|\hat{\gamma}^{\prime}(v_{0})|=\frac{1}{\kappa_{1}(p)^{2}}\sqrt{(\kappa_{1})(p)^{4}G(p)+(\kappa_{1})_{v}(p)^{2}}=\frac{\delta}{\kappa_{1}(p)^{2}},$
we have the conclusion. ∎
By the expression as in Proposition 4.5, when the parallel surface $f^{t}$ with $t=1/\kappa_{1}(p)$ is a swallowtail at $p$, $\kappa_{s}^{C_{1}}$ takes either a positive or negative value at $p$. However, when $f^{t}$ is either a cuspidal lips or a cuspidal beaks at $p$, one can notice the following.
Theorem 4.6.
Let $f\colon U\to\mathbb{R}^{3}$ be a regular surface whose parallel surface $f^{t}$ is a cuspidal lips (resp. a cuspidal beaks) at $p\in U$, where $t=1/\kappa_{1}(p)$. Suppose that $\kappa_{2}(p)=0$. Then the singular curvature $\kappa_{s}^{C_{1}}$ at a cuspidal edge $p$ of $C_{1}$ is positive (resp. negative).
Proof.
Let us take a curvature line coordinate system $(u,v)$. Then by Proposition 3.1, it holds that $(\kappa_{1})_{u}(p)=(\kappa_{1})_{v}(p)=0$ and $\det\operatorname{Hess}(\kappa_{1})(p)>0$ (resp. $\det\operatorname{Hess}(\kappa_{1})(p)<0$) when $f^{t}$ is a cuspidal lips (resp. a cuspidal beaks) at $p$. On the other hand, by Proposition 4.5, the singular curvature $\kappa_{s}^{C_{1}}(p)$ is given by
@@
If the Gaussian curvature $K^{t}$ of $f^{t}$ with $t=1/\kappa_{1}(p)$ is at least ratinally bounded at $p$, then $\kappa_{2}(p)=0$ by Proposition 3.4. Thus we get the following.
Corollary 4.7.
Let $f\colon U\to\mathbb{R}^{3}$ be a regular surface whose parallel surface $f^{t}$ is a cuspidal lips (resp. cuspidal beaks) at $p\in U$, where $t=1/\kappa_{1}(p)$.
  1. (1)
    If the Gaussian curvature $K^{t}$ of $f^{t}$ is rationally bounded at $p$, then the singular curvature $\kappa_{s}^{C_{1}}$ for the focal surface $C_{1}$ of $f$ is positive (resp. negative) at $p$.
  2. (2)
    If the Gaussian curvature $K^{t}$ of $f^{t}$ is bounded near $p$, then $\kappa_{s}^{C_{1}}$ for $C_{1}$ is positive (resp. negative) at $p$.
@@
Moreover, we see that the parallel surface $f^{1}=f+\nu$ is a cuspidal beaks at $(0,0)$ (see Figure 3). We denote the coefficients of the first and second fundamental forms by $E,F,G,L,M$ and $N$. Note that $N=0$. Then one can take a principal vector $\bm{v}_{1}$ with respect to $\kappa_{1}$ is
$\bm{v}_{1}=(\kappa_{1}G,M-\kappa_{1}F).$
This is a null vector of $C_{1}$. It is known that $(\bm{v}_{1}\kappa_{1})^{-1}(0)$ corresponds to the set of singular points of $C_{1}$ ([20]). We remark that $((\bm{v}_{1}\kappa_{1})_{u},(\bm{v}_{1}\kappa_{1})_{v})=(21,6)\neq(0,0)$ at $(0,0)$. Thus a vector field
$\xi=-(\bm{v}_{1}\kappa_{1})_{v}\partial_{u}+(\bm{v}_{1}\kappa_{1})_{u}\partial_{v}$
is tangent to $(\bm{v}_{1}\kappa_{1})^{-1}(0)$. Then it follows that $\det(\xi{C_{1}},\xi\xi{C_{1}},\bm{e}_{1})(0,0)=15876>0$, where $\bm{e}_{1}=\bm{v}_{1}f/|\bm{v}_{1}f|$ is a unit normal vector to $C_{1}$. On the other hand, it holds that
$\bm{v}_{1}\lambda^{C_{1}}=21\quad\det(\xi,\bm{v}_{1})=-21$
at $(0,0)$. Thus we have
$\kappa_{s}^{C_{1}}(0)=\operatorname{sgn}(\bm{v}_{1}\lambda^{C_{1}}(0,0)\cdot\det(\xi,\bm{v}_{1})(0,0))\frac{\det(\xi{C_{1}},\xi\xi{C_{1}},\bm{e}_{1})}{|\xi{C_{1}}|^{3}}(0,0)=-\frac{12}{7}<0.$
@@
$f(u,v)=\left(u,v,\frac{1}{2}u+uv^{2}+u^{4}\right).$
@@
is a unit normal vector to $f$. By direct calculation, we see that $f^{1}=f+\nu$ is a cuspidal lips at $(0,0)$. Moreover, following similar calculations as in the previous example, it can be observed that the singular curvature $\kappa_{s}^{C_{1}}$ of the focal surface $C_{1}$ is positive at $(0,0)$ (see Figure 4). Indeed, we have $\kappa_{s}^{C_{1}}(0)=8>0$.
@@
Appendix A Gaussian curvature of the focal surface
For a surface $f\colon U\to\mathbb{R}^{3}$ parametrized by curvature line coordinates $(u,v)$, we recall the formula for the Gaussian curvature of the focal surface $C_{1}$. By (4.2), the first fundamental form of $C_{1}$ is
(A.1) $\operatorname{I}^{C_{1}}=\frac{(\kappa_{1})_{u}^{2}}{\kappa_{1}^{4}}du^{2}+\frac{2(\kappa_{1})_{u}(\kappa_{1})_{v}}{\kappa_{1}^{4}}dudv+\left(\left(1-\frac{\kappa_{2}}{\kappa_{1}}\right)^{2}G+\frac{(\kappa_{1})_{v}^{2}}{\kappa_{1}^{4}}\right)dv^{2}.$
Thus we get
(A.2) $\det\operatorname{I}^{C_{1}}=\frac{(\kappa_{1})_{u}^{2}}{\kappa_{1}^{4}}\left(1-\frac{\kappa_{2}}{\kappa_{1}}\right)^{2}G=\frac{(\kappa_{1})_{u}^{2}(\kappa_{1}-\kappa_{2})^{2}G}{\kappa_{1}^{6}}.$
On the other hand, the second fundamental form $\operatorname{II}^{C_{1}}$ of $C_{1}$ with respect to $\bm{n}$ is given by
(A.3) $\operatorname{II}^{C_{1}}=-\frac{(\kappa_{1})_{u}L}{\kappa_{1}^{2}|f_{u}|}du^{2}+\frac{G_{u}}{2|f_{u}|}\left(1-\frac{\kappa_{2}}{\kappa_{1}}\right)dv^{2}.$
By Lemma 2.1, we see that
$\frac{G_{u}}{2}=\frac{(\kappa_{2})_{u}G}{\kappa_{1}-\kappa_{2}}.$
Substituting this for (A.3), we obtain
(A.4) $\operatorname{II}^{C_{1}}=-\frac{(\kappa_{1})_{u}L}{\kappa_{1}^{2}|f_{u}|}du^{2}+\frac{(\kappa_{2})_{u}G}{\kappa_{1}|f_{u}|}dv^{2}.$
Thus we have
(A.5) $\det\operatorname{II}^{C_{1}}=-\frac{(\kappa_{1})_{u}(\kappa_{2})_{u}GL}{E\kappa_{1}^{3}}=-\frac{(\kappa_{1})_{u}(\kappa_{2})_{u}G}{\kappa_{1}^{2}}.$
By (A.2) and (A.5), the Gaussian curvature $K^{C_{1}}$ of $C_{1}$ can be calculated as
(A.6) $K^{C_{1}}=\frac{\det\operatorname{II}^{C_{1}}}{\det\operatorname{I}^{C_{1}}}=-\frac{\kappa_{1}^{4}(\kappa_{2})_{u}}{(\kappa_{1})_{u}(\kappa_{1}-\kappa_{2})^{2}}.$
Thus the Gaussian curvature $K^{C_{1}}$ of $C_{1}$ is bounded near a singular point $p$ if and only if $((\kappa_{1})_{u})^{-1}(0)\subset((\kappa_{2})_{u})^{-1}(0)$. In particular, if $K^{C_{1}}$ is bounded and takes a non-zero value at $p$, then $((\kappa_{1})_{u})^{-1}(0)=((\kappa_{2})_{u})^{-1}(0)$.

### v2

@@ 4.1. Geometric invariants at cuspidal edges of focal surfaces
by (2.6). By (2.3), (4.7), (4.10) and Lemma 2.1, we have
$|\hat{\gamma}^{\prime}(v_{0})|^{2}$ $=\frac{1}{\kappa_{1}(p)^{4}}\left((\kappa_{1}(p)-\kappa_{2}(p))^{2}\kappa_{1}(p)^{2}G(p)+(\kappa_{1})_{v}(p)^{2}\right),$
$\left\langle{\hat{\gamma}^{\prime\prime}(v_{0})},{\bm{e}_{1}(p)}\right\rangle$ $=\frac{\kappa_{1}(p)-\kappa_{2}(p)}{\kappa_{1}(p)}\Gamma^{1}_{22}(p)\sqrt{E(p)}=-\frac{\kappa_{1}(p)-\kappa_{2}(p)}{\kappa_{1}(p)}\frac{G_{u}}{2\sqrt{E(p)}}=-\frac{(\kappa_{2})_{u}(p)G(p)}{\kappa_{1}(p)\sqrt{E(p)}}.$
Thus we obtain the assertion for $\kappa_{\nu}^{C_{1}}$.
We next consider $\kappa_{s}^{C_{1}}$. The singular curvature $\kappa_{s}^{C_{1}}$ at $p$ is
$\kappa_{s}^{C_{1}}(p)=\operatorname{sgn}(\det(\gamma^{\prime},\eta^{C_{1}})(v_{0})\cdot\eta^{C_{1}}\lambda^{C_{1}}(p))\frac{\det(\hat{\gamma}^{\prime}(v_{0}),\hat{\gamma}^{\prime\prime}(v_{0}),\bm{e}_{1}(p))}{|\hat{\gamma}^{\prime}(v_{0})|^{3}}.$
@@
$\bm{e}_{1}(p)\times\hat{\gamma}^{\prime}(v_{0})=\frac{(\kappa_{1})_{v}(p)}{\kappa_{1}(p)^{2}}\bm{e}_{2}(p)+\frac{\kappa_{1}(p)-\kappa_{2}(p)}{\kappa_{1}(p)}\sqrt{G(p)}\nu(p).$
@@
$\det(\hat{\gamma}^{\prime}(v_{0}),\hat{\gamma}^{\prime\prime}(v_{0}),\bm{e}_{1}(p))=\left\langle{\bm{e}_{1}(p)\times\hat{\gamma}^{\prime}(v_{0})},{\hat{\gamma}^{\prime\prime}(v_{0})}\right\rangle=\frac{\sqrt{G}(p)}{\kappa_{1}(p)^{3}}X_{1}(p).$
@@
$\eta^{C_{1}}\lambda^{C_{1}}(p)=\frac{(\kappa_{1})_{uu}(p)(\kappa_{1}(p)-\kappa_{2}(p))}{\kappa_{1}(p)^{3}}|f_{v}(p)|,\quad\det(\gamma^{\prime},\eta)(v_{0})=-1$
@@
$\operatorname{sgn}(\det(\gamma^{\prime},\eta)(v_{0})\cdot\eta^{C_{1}}\lambda^{C_{1}}(p))=-\operatorname{sgn}((\kappa_{1})_{uu}(p)\kappa_{1}(p)(\kappa_{1}(p)-\kappa_{2}(p))).$
Thus we have the conclusion for $\kappa_{s}^{C_{1}}$.
For $C_{2}$, one can show in a similar manner. ∎
By using Proposition 4.4, we show Theorem 1.2.
Proof of Theorem 1.2.
We show the case of $C_{1}$. By Proposition 1.1, $\kappa_{2}(p)=0$. Let us take a curvature line coordinate system $(u,v)$. Then by Proposition 3.1, it holds that $(\kappa_{1})_{u}(p)=(\kappa_{1})_{v}(p)=0$ and $\det\operatorname{Hess}(\kappa_{1})(p)>0$ (resp. $\det\operatorname{Hess}(\kappa_{1})(p)<0$) when $f^{t}$ is a cuspidal lips (resp. a cuspidal beaks) at $p$. On the other hand, by Proposition 4.4, the singular curvature $\kappa_{s}^{C_{1}}(p)$ is given by
@@
Moreover, we see that the parallel surface $f^{1}=f+\nu$ is a cuspidal beaks at $(0,0)$ (see Figure 3). We find that the limiting normal curvature of $f^{1}$ vanishes at $(0,0)$. In addition, the singular curvature $\kappa_{s}^{C_{1}}$ of $C_{1}$ is $\kappa_{s}^{C_{1}}(0)=-12/7<0$.
@@
$f(u,v)=\left(u,v,\frac{1}{2}u^{2}+uv^{2}+u^{4}\right).$
@@
is a unit normal vector to $f$. By direct calculation, we see that $f^{1}=f+\nu$ is a cuspidal lips at $(0,0)$. Moreover, it can be observed that the singular curvature $\kappa_{s}^{C_{1}}$ of the focal surface $C_{1}$ is $\kappa_{s}^{C_{1}}(0)=8>0$ at $(0,0)$ (see Figure 4).
@@
Example 4.7.
Let $f\colon\mathbb{R}^{2}\to\mathbb{R}^{3}$ be a regular surface given by
$f(u,v)=\left(u,v,\frac{1}{2}u^{2}-v^{2}+u^{4}+u^{3}v\right).$
Then the unit normal vector field $\nu$ is
$\nu(u,v)=\frac{(-u(1+4u^{2}+3uv),-u^{3}+2v,1)}{\sqrt{1+(u^{3}+2v)^{2}+u^{2}(1+4u^{2}+3uv)^{2}}}.$
Principal curvatures $\kappa_{1},\kappa_{2}$ of $f$ can be chosen as
$\kappa_{1}(0,0)=-2,\quad\kappa_{2}(0,0)=1\neq 0.$
Then a parallel surface $f^{-1/2}=f-\nu/2$ has a cuspidal lips at $(0,0)$. Moreover, the singular curvature of $C_{1}$ at the origin is calculated as $\kappa_{s}^{C_{1}}(0)=-9/4<0$ (see Figure 5) This gives a counterexample of Theorem 1.2.
Figure 5. The image of $f^{-1/2}$ (left), $C_{1}$ (middle) and both of them (right).
If the Gaussian curvature of $f^{t}$ is bounded near its singular point $p$, then $\kappa_{\nu}^{t}$ vanishes at $p$ (see Appendix A). Thus we get the following by Theorem 1.2 immediately.
Corollary 4.8.
Let $f\colon U\to\mathbb{R}^{3}$ be a regular surface whose parallel surface $f^{t}$ is a cuspidal lips (resp. cuspidal beaks) at $p\in U$, where $t=1/\kappa_{i}(p)$ ($i=1$ or $2$). If the Gaussian curvature $K^{t}$ of $f^{t}$ is bounded near $p$, then $\kappa_{s}^{C_{i}}$ for $C_{i}$ is positive (resp. negative) at $p$.
Appendix A Gaussian curvature of parallels and focal surfaces
Let $f\colon U\to\mathbb{R}^{3}$ be a regular surface and $\nu$ its unit normal vector field of $f$. Then the Gaussian curvature $K^{t}$ of the parallel surface $f^{t}=f+t\nu$ is
$K^{t}=\frac{\kappa_{1}\kappa_{2}}{(1-t\kappa_{1})(1-t\kappa_{2})}$
(cf. [4]). Thus when $t=1/\kappa_{1}(p)$ (resp. $t=1/\kappa_{2}(p)$), $K^{t}$ is bounded near $p$ if and only if $\kappa_{1}^{-1}(\kappa_{1}(p))\subset\kappa_{2}^{-1}(0)$ (resp. $\kappa_{2}^{-1}(\kappa_{2}(p))\subset\kappa_{1}^{-1}(0)$) (cf. [20, 12]). Moreover, the Gaussian curvature $K^{C_{i}}$ ($i=1,2$) of $C_{i}$ can be calculated as
(A.1) $K^{C_{i}}=\begin{dcases}-\frac{\kappa_{1}^{4}(\kappa_{2})_{u}}{(\kappa_{1})_{u}(\kappa_{1}-\kappa_{2})^{2}}&(i=1),\\ -\frac{\kappa_{2}^{4}(\kappa_{1})_{v}}{(\kappa_{2})_{v}(\kappa_{1}-\kappa_{2})^{2}}&(i=2)\end{dcases}$
([15, 4]). Thus the Gaussian curvature $K^{C_{1}}$ of $C_{1}$ is bounded near a singular point $p$ if and only if $((\kappa_{1})_{u})^{-1}(0)\subset((\kappa_{2})_{u})^{-1}(0)$. In particular, if $K^{C_{1}}$ is bounded and takes a non-zero value at $p$, then $((\kappa_{1})_{u})^{-1}(0)=((\kappa_{2})_{u})^{-1}(0)$.